If you react .994 grams of iron (III) oxide with aluminum, what mass of iron (Fe) will be produced?

                                Fe2O3   +   Al   ®   Al2O3  +  F


1
Expert's answer
2021-03-01T14:56:31-0500
"m_{Fe}=2m_{Fe_2O_3}\frac{M_{Fe}}{M_{Fe_2O_3}}\\
m_{Fe}=2(0.994)\frac{55.845}{159.69}=0.695\ g"